# Subtleties in the pseudomodes formalism

### v1

@@ Appendix C Non-Diagonalizable Configuration as the Limit of a Diagonalizable One
$a_{1}$ $=\frac{1}{2}\left(\zeta_{1}^{2}+\zeta_{2}^{2}+\frac{2\zeta_{1}\zeta_{2}(\epsilon+1)}{\sqrt{\epsilon(\epsilon+2)}}\right)\,$, (163)
$a_{2}$ $=\frac{1}{2}\left(\zeta_{1}^{2}+\zeta_{2}^{2}-\frac{2\zeta_{2}\zeta_{1}(\epsilon+1)}{\sqrt{\epsilon(\epsilon+2)}}\right)\,$, (164)
$b_{1}$ $=\frac{\zeta_{1}^{2}-\zeta_{2}^{2}}{2\sqrt{\epsilon(\epsilon+2)}}\,$, (165)
$b_{2}$ $=\frac{\zeta_{2}^{2}-\zeta_{1}^{2}}{2\sqrt{\epsilon(\epsilon+2)}}\,.$ (166)
Note that each of these diverges as $\epsilon\to 0$. However, if we construct the spectral density by
$J(\omega)=\sum_{i=1}^{n_{\alpha}}a_{i}\frac{\gamma_{i}}{(\omega-\varepsilon_{i})^{2}+(\gamma_{i}/2)^{2}}-b_{i}\frac{2(\omega-\varepsilon_{i})}{(\omega-\varepsilon_{i})^{2}+(\gamma_{i}/2)^{2}}\,$, (167)
where $\epsilon_{i},\gamma_{i}$ come from the eigenvalues of $W$, then we obtain:
Taking the limit of $\epsilon\to 0$, we obtain:
exactly as before.
Appendix D Prony’s Method
Suppose we have a memory kernel $\chi(t)$ that we want to approximate as:
$\chi(t)=\sum_{k=1}^{n}\kappa_{k}e^{-\lambda_{k}t}\,$, (170)
for complex parameters $\kappa_{k},\lambda_{k}$, where we want $n$ as small as possible to get a good fit. Prony’s method Wilson et al. (2019); Almunif et al. (2020); Chen et al. (2022) consists of the following steps:
  1. 1.
    Choose a time interval $\Delta t$ and cutoff time $t_{c}=2N\Delta t$ to sample the correlation function, where $N\geq 2L+1$.
  2. 2.
    Let $\phi_{j}=C(j\Delta t)$ for $j=0,\dots,2N$ and construct the $(N+1)\times(N+1)$ matrix $\mathbf{H}_{ij}=\phi_{i+j+1}$.
  3. 3.
    Compute the Takagi decomposition Ikramov (2012): find vectors $\mathbf{u}_{m}$ such that $\mathbf{H}\mathbf{u}_{m}=\sigma_{m}\mathbf{u}_{m}^{*}$. This is analogous to finding the eigendecomposition of a Hermitian matrix as $\mathbf{H}{|{m}\rangle}=\sigma_{m}{|{m}\rangle}$, but for a complex-symmetric matrix. The matrix $U$ whose columns are $\mathbf{u}_{m}$ is unitary so we could alternately write $H=U\sigma U^{T}$ where we take the transpose of $U$ rather than the adjoint. We order the $\sigma_{m}$ by descending values of $|\sigma_{m}|$ and the $\mathbf{u}_{m}$ accordingly.
  4. 4.
    Define the polynomial $f(z)=\sum_{n=0}^{N}\left(\mathbf{u}_{L}\right)_{n}z^{n}$ and obtain its roots $w_{0},w_{1},\dots w_{N}$ ordered by increasing modulus.
  5. 5.
    Of these roots, the first $L$ will have $|w_{k}|<1$. Write these roots as $r_{k}e^{i\theta_{k}}$ with $\pi<\theta\leq\pi$ and choose the exponents to be $\lambda_{k}=-\frac{2N}{t_{c}}\left[\log(r_{k})+i\theta_{k}\right]$.
  6. 6.
    Obtain $\kappa_{k}$ by least-squares fitting the equations $\phi_{j}=\sum_{k=1}^{L}\kappa_{k}w_{k}^{j}$ for $j=0,1,\dots 2N$.
Fitting with Prony’s method can be quite sensitive to the choice made in step 1. Thus, it is generally most practical to sample a large number of interval choices and then select the choice that gives the effective spectral density that best fits the true spectral density, or to perform optimization on the endpoints of the time interval used.
Appendix E 3x3 non-diagonalizable Configuration
We compute the eigenvalues of this matrix as:
$x=\sqrt{\frac{1}{2}\left(5\sqrt{5}-11\right)}\,$, (175)
and so we find a candidate 3x3 non-diagonalizable $W$ as:
and effective spectral density
for this pseudomode configuration.

### v2

@@ Appendix F Fit Parameters in Figures 2 and 4
[…]
0.962 -0.136 - 0.0644ii 0.0696+0.122ii -0.227+0.258ii 0.395+0.404ii 0.000761 0.0167-0.000388ii 0.148-0.0934ii
1.383 0.0132 + 0.169ii 0.164-0.237ii 0.239+0.250ii -0.255+0.434ii 0.0167+0.000388ii 0.0140 -0.481-0.485ii
6.575 -0.103 + 0.117ii 0.0427+0.00337ii 0.131+0.253ii 0.337-0.174ii 0.148+0.0934ii -0.481+0.485ii -0.00211
Table 4: Fit parameters for the diagonalizable fit used in Fig. 2(a,c), as they appear in (27).
$\kappa_{k}$ $\varepsilon_{k}$ $\gamma_{k}$
0.67900 + 0.37224ii -0.34457 -1.23329
0.67900 - 0.37224ii 0.34457 -1.23329
0.10729 - 0.32214ii 0.79313 -0.62406
0.10729 + 0.32214ii -0.79313 -0.62406
-0.00089 - 0.07379ii 0.95949 -0.15409
-0.00089 + 0.07379ii -0.95949 -0.15409
Table 5: Fit parameters for the uncoupled modes fit used in Fig. 4, as they appear in (20).
$\kappa_{k}$ $\varepsilon_{k}$ $\gamma_{k}$
0.0369 -0.706 0.295
0.0369 0.706 0.295
0.0699 -0.377 0.423
0.0699 0.377 0.423
[…]